# Oscillations Make Neural Networks Robust to Quantization

### v1

@@ Abstract
We challenge the prevailing view that oscillations in Quantization Aware Training (QAT) are merely undesirable artifacts caused by the Straight-Through Estimator (STE). Through theoretical analysis of QAT in linear models, we demonstrate that the gradient of the loss function can be decomposed into two terms: the original full-precision loss and a term that causes quantization oscillations. Based on these insights, we propose a novel regularization method that induces oscillations to improve quantization robustness. Contrary to traditional methods that focuses on minimizing the effects of oscillations, our approach leverages the beneficial aspects of weight oscillations to preserve model performance under quantization. Our empirical results on ResNet-18 and Tiny ViT demonstrate that this counter-intuitive strategy matches QAT accuracy at $\geq$ 3-bit weight quantization, while maintaining close to full precision accuracy at bits greater than the target bit. Our work therefore provides a new perspective on model preparation for quantization, particularly for finding weights that are robust to changes in the bit of the quantizer – an area where current methods struggle to match the accuracy of QAT at specific bits.  [^1] Source code available at: https://github.com/saintslab/OsciQuant
Section 1 Introduction
Quantization is the mapping of continuous values to discrete values. In neural networks, quantization reduces the computational complexity and memory requirements by representing weights and/or activations with fewer bits Gupta et al. (2015). In the case of weight only quantization, this means applying a quantizer $q(\cdot)$ to the network’s weights $\mathbf{w}$, with an additional implicit goal of maintaining the original performance i.e. $\mathcal{L}(q(\mathbf{w}))\approx\mathcal{L}(\mathbf{w})$, where $\mathcal{L}(\cdot)$ is a loss function.
When training neural networks intended for quantization, an essential step during optimization is accounting for the effects of applying a quantizer to the weights. Quantization introduces a perturbation to the weights. For uniform quantizers, this is bounded by $\frac{s}{2}$, where $s$ is the scale factor. At higher bit widths ($\geq 8$ bits), this perturbation is small, and standard training procedures often yield weights that are resilient to quantization noise Nagel et al. (2021). In such cases, applying quantization after training, known as Post-Training Quantization (PTQ), is sufficient to maintain acceptable performance levels Nagel et al. (2021).
However, as we reduce the bit width to lower precision ($\leq 4$ bits), the quantization perturbation becomes more significant, and the model’s performance tends to degrade substantially after quantization. This is because the increased perturbation can lead to larger discrepancy between $q(\mathbf{w})$ and $\mathbf{w}$. To address this challenge, much research has gone into finding strategies to mitigate the effects of quantization on model accuracy, ensuring that the network remains accurate even after low-bit quantization.
Figure 1: Oscillatory behavior in Quantization-Aware Training (QAT) for a simple linear model. The figure shows a quantized linear model $f(x)=q(w)x$ with a single weight $w$, where $x=1$ and target output $y=0.75$. When doing squared loss with QAT an additional term is introduced to the gradient (Eq. 14), which causes w to oscillate around the quantization threshold. This oscillation results in $q(w)$ alternating between the 0 and 1 quantization bins.
Though many methods have been proposed for mitigating the accuracy degradation due to quantization, Quantization-Aware Training (QAT) Jacob et al. (2018) remains one of the most widely adopted approaches. QAT works by incorporating quantization effects directly into the training process - quantizing weights in the forward pass while using the Straight-Through Estimator (STE) Bengio et al. (2013) for gradient approximation during backpropagation. Research has identified an interesting phenomenon in QAT with the STE, known as weight oscillations, where the quantized weights alternate between two adjacent quantized states during training Défossez et al. (2021); Nagel et al. (2022). While traditionally viewed as a detrimental effect that should be suppressed through dampening or weight freezing techniques, there also exists evidence suggesting these oscillations might play a more nuanced role in the training dynamics of QAT.
We claim that weight oscillations during training are beneficial and that indeed they are the driving mechanism behind QAT. Our primary contributions that support this claim are:
@@
    we isolate the mechanism that leads to weight oscillations during QAT (Sec. 4);
@@
    we develop a regularization method that induces weight oscillations during training using this mechanism (Sec. 5);
@@
    we show experimentally that weight oscillations are sufficient for preserving performance after quantization on small-scale computer vision tasks (Sec. 6).
Since previous results have shown that weights oscillations are also necessary for good quantization performance with QAT (see Sec. 7 for details), and extrapolating from our experiments, our results suggest that weight oscillations capture all the beneficial effects of QAT while avoiding unintended side-effects. For instance, in our experiments our method avoids overfitting to the bit-width used during training, resulting in superior cross-quantization performance compared to QAT.
Section 2 Related Work
The most used strategy to minimize the impact of quantization on model accuracy is to minimize the quantization error. This can be achieved by adjusting the granularity of the quantizer—for instance, using per-channel Nagel et al. (2019) or block-wise quantization Dettmers et al. (2022) instead of per-tensor quantization. While these methods reduce quantization error without additional training, they come with increased storage requirements due to extra quantization parameters and may still fall short at very low bit widths, necessitating the combination with other approaches.
Consequently, extensive research has been dedicated to developing techniques that explicitly minimize the quantization error during optimization Hung et al. (2015); Hirose et al. (2017); Li et al. (2019); Choi et al. (2020); Han et al. (2021); Zhong et al. (2025). Given a model $\mathbf{w}$ the hope is that by ensuring $q(\mathbf{w})\approx\mathbf{w}$, we likely also have $\mathcal{L}(q(\mathbf{w}))\approx\mathcal{L}(\mathbf{w})$, thereby preserving model accuracy after quantization.
An alternative and less explored approach involves training models to be robust to quantization perturbations without necessarily minimizing the quantization error itself. This means finding weights $\mathbf{w}$ such that $\mathcal{L}(q(\mathbf{w}))\approx\mathcal{L}(\mathbf{w})$ even if $q(\mathbf{w})$ is not close to $\mathbf{w}$ Alizadeh et al. (2020); Chmiel et al. (2020). Such methods focus on enhancing the robustness of the model to the quantization error, leading to better performance at bits different than the ones used in the quantizer, which we will refer to as cross-bit quantization.
A third approach is to train supernets on the desired configurations of the quantizers Xu et al. ; Xu et al. (2023). This approach increases the training complexity and cost, which is not incurred by explicit regularization.
Despite these efforts, the aforementioned strategies often fall short of the accuracy obtained with QAT Jacob et al. (2018) at individual bits or indirectly rely upon QAT themselves. In short, QAT integrates the quantization process into the training loop allowing the model to adapt to the quantization effects directly. This is done by quantizing the weights during the forward pass and using techniques like the Straight-Through-Estimator (STE) to approximate the gradient of the quantizer (Which has a derivative of zero almost everywhere) during backpropagation Bengio et al. (2013).
Yet, there is limited understanding of how QAT affects model optimization and why it outperforms other methods. One phenomenon observed during QAT is weight oscillations Défossez et al. (2021); Nagel et al. (2022), which are periodic changes in the value of the quantized weight between two adjacent quantization levels. It is speculated in these works that that the abrupt changes in values caused by oscillations can interfere negatively with optimization. Oscillations are assumed to be undesirable side effects caused by the use of the STE during backpropagation, as the STE allows gradients to pass through the rounding operation in the quantizer, which has a gradient of zero almost everywhere Défossez et al. (2021); Nagel et al. (2022).
Several approaches have been suggested to mitigate oscillations, such as dampening or freezing the oscillating weights, which have shown improved accuracy Nagel et al. (2022); Gupta and Asthana (2024). However, the reported gains are sometimes marginal, and these methods may inadvertently also hinder the optimization process. For instance, Nagel et al. (2022) notes that freezing or dampening weights too early during training can hurt optimization, indicating that oscillations might contribute to finding better quantized minima. Liu et al. (2023) propose that weights with low oscillation frequency should be frozen, where as high-frequency ones should be left unfrozen, under the rational that high frequency means the network has little confidence in what value to quantize the weight to, where as low frequency means the optimal weight lies close to a quantization level.
Though QAT often provides the best accuracy for a given target bit, degradation to a lesser or greater extent exists when the bit of the quantizer is different to the one seen during training, ie. cross-bit quantization Alizadeh et al. (2020); Chmiel et al. (2020). This means QAT requires training and storing of weights for each desired bit width. This specialization can also pose challenges when deploying models across different hardware platforms, each potentially using different quantization schemes Reddi et al. (2020), making it difficult to develop models which can be easily quantized at deployment according to end-user requirements.
This makes robust quantization methods an interesting research avenue, especially if they could be improved to match the individual bit performance of QAT. In this work, we aim to deepen the understanding of how QAT influences model optimization, particularly focusing on the role of weight oscillations and their relation to robustness.
Section 3 Preliminaries
Section 3.1 Quantization
@@
We consider a uniform symmetric quantizer with a max-range scale factor. The quantization operation $q(\cdot)$ can then be expressed as
$q(\mathbf{w})$ $=s\cdot\left\lceil\frac{\mathbf{w}}{s}\right\rfloor$ (1)
[…]
The scale factor $s$ is set to cover the range of $\mathbf{w}$ as this removes the need for the usual clamping operation in the quantizer, while keeping the number of bins symmetric around 0:
$s$ $=\frac{\max(\lvert\alpha\rvert,\lvert\beta\rvert)}{2^{b-1}-1}$ (2)
Where $b$ is the bit in the quantizer and $\alpha,\beta$ are the min. and max. values respectively of the layer wise weight $\mathbf{w}$.
@@
$\mathbf{\Delta}(\mathbf{w})$ $=\mathbf{w}-q(\mathbf{w})$ (3)
@@
Due to the uniform quantizer, for all bins the absolute error is bounded between $20\leq|\mathbf{\Delta}|\leq s/$ 2, which is maximized at quantization thresholds and 0 at quantization levels.
Section 3.2 Quantization-Aware Training
@@
While there exist many variants of QAT, fundamentally the forward pass is performed using the quantized weights $q(\mathbf{w})$ in most variants of QAT Jacob et al. (2018); Krishnamoorthi (2018), simulating the effect of using low-precision weights. In principle the gradient for the weights during QAT is given by:
$\frac{\partial\mathcal{L}(q(\mathbf{w}))}{\partial\mathbf{w}}=\frac{\partial\mathcal{L}(q(\mathbf{w}))}{\partial q(\mathbf{w})}\cdot\frac{\partial q(\mathbf{w})}{\partial\mathbf{w}}$ (4)
@@
A problem with the above formulation is that the gradient of the rounding operation used in the quantizer is zero almost everywhere, causing the last term to interrupt gradient-based learning. A popular solution to this problem is to use the so-called Straight-Through Estimator (STE) Bengio et al. (2013). We define the STE to be the operator ^^$\frac{\hat{\partial}}{\hat{\partial}\mathbf{x}}$ such that ^$\frac{\hat{\partial}f}{\hat{\partial}x}$ is obtained by computing $\frac{\partial f}{\partial x}$ and in the resulting expression replacing $q^{\prime}$ (the derivative of $q$) by the constant function equal to $1$. In other words, if $\frac{\partial f}{\partial x}=g(\ldots,q^{\prime},\ldots)$ then ^$\frac{\hat{\partial}f}{\hat{\partial}x}=g(\ldots,1,\ldots)$.
Section 4 Oscillations in QAT
@@
Previous studies have explored linear models to analyze the behavior of QAT and the phenomenon of weight oscillations Défossez et al. (2021); Nagel et al. (2022); Liu et al. (2023); Gupta and Asthana (2024). Inspired by these works, we analyze a linear regression model to gain theoretical insights into the optimization dynamics during QAT.
@@
Consider a linear model with a single weight $w$, input $x$ and target $y$ $\in\mathbb{R}$. The quantized version of this model is defined as $f(x)=q(w)x$, where $q(\cdot)$ is the quantizer from Eq. 1. The quadratic loss for the quantized model is given by
$\mathcal{L}(q(w))=\frac{1}{2}(q(w)x-y)^{2}.$ (5)
@@
Our goal in this section is to understand how QAT affects the full precision optimization process. For a given loss function $\mathcal{L}(\cdot)$ with quantized weights, we have
$\mathcal{L}(q(w))$ $=\mathcal{L}(w)+\mathcal{L}(q(w))-\mathcal{L}(w)$ (6)
We can then expand the difference caused by quantization as follows
$\delta_{\text{$\mathcal{L}$}}$ $=\mathcal{L}(q(w))-\mathcal{L}(w)$ (7)
$=\frac{1}{2}\left((q(w)x-y)^{2}-(wx-y)^{2}\right)$ (8)
$=\frac{1}{2}\left((q(w)x)^{2}-(wx)^{2}-2y(q(w)x-wx)\right)$ (9)
$=\frac{1}{2}\left(x^{2}\left(q(w)^{2}-w^{2}\right)\right)+\left(yx(w-q(w))\right)$ (10)
@@
This expression decomposes the loss difference into a quadratic term $\frac{1}{2}x^{2}(q(w)^{2}-w^{2})$ and a linear term $yx(w-q(w))$.
Next we derive the gradient of $\delta_{\mathcal{L}}$ wrt. $w$:
$\frac{\partial\delta_{\mathcal{L}}}{\partial w}$ $=\frac{\partial}{\partial w}\bigg{(}\mathcal{L}(q(w))-\mathcal{L}(w)\bigg{)}$ (11)
$=\frac{\partial}{\partial w}\left(\frac{1}{2}x^{2}(q(w)^{2}-w^{2})+yx(w-q(w))\right)$ (12)
$=x^{2}\left(q(w)\frac{\partial q(w)}{\partial w}-w\right)+yx\left(1-\frac{\partial q(w)}{\partial w}\right)$ (13)
@@
Using the STE and recalling that ^$\frac{\hat{\partial}q}{\hat{\partial}w}=1$ the expression of the STE gradient simplifies to
^$\frac{\hat{\partial}\delta_{\mathcal{L}}}{\hat{\partial}w}=x^{2}(q(w)-w)=-x^{2}\mathbf{\Delta}(w).$ (14)

### v2

@@ Abstract
We challenge the prevailing view that weight oscillations observed during Quantization Aware Training (QAT) are merely undesirable side-effects and argue instead that they are an essential part of QAT. We show in a univariate linear model that QAT results in an additional loss term that causes oscillations by pushing weights away from their nearest quantization level. Based on the mechanism from the analysis, we then derive a regularizer that induces oscillations in the weights of neural networks during training. Our empirical results on ResNet-18 and Tiny Vision Transformer, evaluated on CIFAR-10 and Tiny ImageNet datasets, demonstrate across a range of quantization levels that training with oscillations followed by post-training quantization (PTQ) is sufficient to recover the performance of QAT in most cases. With this work we provide further insight into the dynamics of QAT and contribute a novel insight into explaining the role of oscillations in QAT which until now have been considered to have a primarily negative effect on quantization. [^1] Source code is available at https://github.com/saintslab/osc_reg.
1 Introduction
Deep neural networks have grown increasingly successful at solving difficult modelling tasks, but are also increasingly becoming more expensive to use. As model sizes balloon into the hundreds of millions of parameters, the cost of inference has become a significant bottleneck, particularly for deployment on edge devices or usage in large-scale services sevilla2022compute. To reduce this cost, quantization of model weights has emerged as a prominent strategy nagel2021white.
Weight quantization, however, comes with its own cost, namely a drop in accuracy. Reducing precision introduces quantization error, and naive approaches like Post-Training Quantization (PTQ) often lead to sharp degradations in model performance at low bit widths. To combat this degradation in accuracy, many methods have been proposed, usually centered around minimizing the quantization error BridgeDeepLearning; minimizeQuantError; APTQ; Choi_2020; ImprovingLowBit; zhong2024mbquantnovelmultibranchtopology.
Figure 1: Oscillatory behavior during QAT in a one weight linear model $\hat{y}=q(w)x$, with weight $w$, quantized weight $q(w)$, input $x=1$ and target $y=0.75$. The gradient of the loss term of this model during QAT can be decomposed into two terms: $\nabla L(w)$ and $\nabla\delta_{L}=q(w)-w$, where the latter term is what differentiates QAT from just optimizing the full-precision loss $L(w)$. During QAT, $\nabla L(w)$ always points towards $y$, while $\nabla\delta_{L}$ introduces a dynamic which pushes $w$ towards the nearest bin threshold. This causes $w$ to oscillate when $y$ is not exactly on a quantization level. In the above case this makes $q(w)$ alternate between 0 and 1. Note the frequency of oscillations of $q(w)$ lets the quantized weight on average to converge to 0.75.
Yet these approaches fall short of Quantization-Aware Training (QAT) jacob2017quantization; krishnamoorthi2018quantizing, which integrates quantization directly into the training loop, yielding models that remain robust even under low-precision constraints. However, despite its empirical success, the underlying behavior of QAT remains poorly understood. In particular, QAT often exhibits oscillations in the weights during optimization, while regular training does not. Rather than settling into a stable state, quantized weights fluctuate between adjacent quantization levels. These oscillations are widely seen as undesirable artifacts. Several works have tried to suppress these weight oscillations mainly through dampening or weight freezing pseudoQuantNoise; nagel2022overcoming; huang2023quantization; gupta2023reducing; vitoscillations.
This paper poses the question: Have we misunderstood the nature of oscillations in QAT? To the best of our knowledge, we are the first to argue that weight oscillations in QAT are not a flaw, but rather an essential feature of the training dynamics. Using a toy model, we show that oscillations arise from a gradient component in QAT which pushes weights towards the nearest quantization threshold – this stands in contrast to aligning weights at the quantization level, which would minimize the quantization error. We then induce such oscillations to push the weights towards the nearest thresholds deliberately during the model training. This is realized as a simple regularization term that pushes weights towards their nearest quantization threshold. Surprisingly, this leads to models that are robust to quantization and in most cases we find that these regularization-induced oscillations recover the accuracy obtained with QAT. Additionally, in many cases, oscillations outperform QAT under cross-bit evaluation (i.e. testing at precision levels not simulated during training). These findings further deepen our understanding of QAT and point to a more nuanced role for weight oscillations; suggesting they may be beneficial to QAT rather than being solely detrimental as argued in most existing literature.
Our primary contributions supporting this claim are:
@@ 1 Introduction
    Using a univariate model we illustrate how the Straight Through Estimator (STE) leads to oscillations and clustering of model weights during QAT (Sec. 3);
@@
    We show experimentally that by using a mechanism inspired by the toy model, we can induce oscillations and clustering during training in neural networks (Sec. 4).
@@
    We empirically confirm using the CIFAR-10 krizhevsky2009learning and Tiny-ImageNet Le2015TinyIV datasets on a multi-layer perceptron (MLP), ResNet-18 he2016deep and Tiny Vision transformer (Tiny ViT) wu2022tinyvit that introducing oscillations through regularization in most cases recovers the accuracy of QAT (Sec. 5).
2 Preliminaries and Related Work
Quantization:
@@ 2 Preliminaries and Related Work
We consider a uniform symmetric, scalar quantizer, $q(\cdot)$, that can then be expressed as:
$q(\mathbf{w})$ $=s\cdot\left\lceil\frac{\mathbf{w}}{s}\right\rfloor.$ (1)
Here, $s$ represents the scale factor and $\left\lceil\cdot\right\rfloor$ denotes the rounding operation.
The scale factor $s$ is set to cover the range of $\mathbf{w}$ as this removes the need for the clamping operation $\text{clamp}(q(\textbf{w});\alpha,\beta)$ in the quantizer. The function clamp(.) restricts $\alpha\leq q(w_{i})\leq\beta$ and any values less or greater than is set to $\alpha$ and $\beta$ respectively. Additionally we set the number of positive and negative quantization levels to be $2^{b-1}-1$, so we have symmetric number of levels around 0:
$s$ $=\frac{\max(\lvert\textbf{w}\rvert)}{2^{b-1}-1},$ (2)
where $b$ is the number of bits the quantizer can use.
@@
$\mathbf{\Delta}(\mathbf{w})$ $=\mathbf{w}-q(\mathbf{w}).$ (3)
@@
Due to the uniform quantizer each component of the absolute error (due to quantizing $w_{i}$) is bounded between $0\leq|\mathbf{\Delta}_{i}|\leq s/2$, for all the bins. This is maximized at quantization thresholds and is 0 at quantization levels.
Straight-Through Estimator (STE):
STE is a heuristic gradient approximation technique commonly used to enable backpropagation through functions $f(\cdot)$ whose gradients are zero almost everywhere bengio2013ste. In the forward pass, the function $f(\mathbf{x})$ is applied as usual. During the backward pass, however, its gradient is approximated by replacing the derivative of $f$ with the constant $1$. Formally, STE defines the approximate gradient of a function $f$ as
$\frac{{\partial}f}{{\partial}\textbf{x}}\approx\frac{\hat{\partial}f}{\hat{\partial}\textbf{x}}=1.$ (4)
Quantization-Aware Training (QAT):
@@
In QAT, the network maintains full-precision weights for learning but quantizes them during the forward pass before computation. This allows training to proceed with gradients applied to full-precision weights, while the network effectively operates on quantized values. While there exist many variants of QAT, the typical forward pass is performed using the quantized weights $q(\mathbf{w})$ jacob2017quantization; krishnamoorthi2018quantizing. The gradient for the weights during QAT is given by
$\frac{\partial\mathcal{L}(q(\mathbf{w}))}{\partial\mathbf{w}}=\frac{\partial\mathcal{L}(q(\mathbf{w}))}{\partial q(\mathbf{w})}\cdot\frac{\partial q(\mathbf{w})}{\partial\mathbf{w}}.$ (5)
@@
A problem with the above formulation is that the gradient of the quantizer $\frac{\partial q(\mathbf{w})}{\partial\mathbf{w}}$ is zero almost everywhere, causing the last term to interrupt gradient-based learning. A widely used solution to this problem is to use the STE (Eq. (4)).
Weight Oscillations:
We use the definition from nagel2022overcoming which defines a weight oscillation during QAT to occur in iteration $t$ if it satisfies both the following conditions:
  1. 1.
    The quantized value of the weight needs to change between iterations i.e. $q(w_{t})\neq q(w_{t-1})$.
  2. 2.
    The direction of the change in the quantized domain needs to be the opposite than that of its previous change i.e. $\mathrm{sign}(\Gamma_{t})\neq\mathrm{sign}(\Gamma_{\tau})$, where $\tau$ is the iteration of the last change, and $\Gamma_{t}=q(w_{t})-q(w_{t-1})$ is the direction of the change.
Related Work:
Minimizing the quantization error is the most commonly used strategy to reduce the impact of quantization on model accuracy. Extensive research has been dedicated to developing techniques that explicitly minimize the quantization error during optimization BridgeDeepLearning; minimizeQuantError; APTQ; Choi_2020; ImprovingLowBit; zhong2024mbquantnovelmultibranchtopology. Despite these efforts, the aforementioned strategies often fall short of the accuracy obtained with QAT jacob2017quantization at individual bits or indirectly rely upon QAT themselves.
However, there is limited understanding of how QAT affects model optimization and why it outperforms other methods. One phenomenon observed during QAT is weight oscillations pseudoQuantNoise; nagel2022overcoming, which are periodic changes in the value of the quantized weight between two adjacent quantization levels. It is speculated in these works that the abrupt changes in values caused by oscillations can interfere negatively with optimization. Oscillations are assumed to be undesirable side effects caused by the use of the STE during backpropagation, as the STE allows gradients to pass through the rounding operation in the quantizer, which has a gradient of zero almost everywhere pseudoQuantNoise; nagel2022overcoming.
Several approaches have been suggested to mitigate oscillations by either freezing or dampening pseudoQuantNoise; nagel2022overcoming; huang2023quantization; gupta2023reducing; vitoscillations. However, the reported accuracy gains are sometimes marginal, and these methods may inadvertently also hinder the optimization process. For instance, nagel2022overcoming notes that freezing or dampening weights too early during training can hurt optimization, indicating that oscillations might contribute to finding better quantized minima of the loss. vitoscillations propose that weights with low oscillation frequency should be frozen, where as high-frequency ones should be left unfrozen, under the rationale that high frequency means the network has little confidence in what value to quantize the weight to, whereas low frequency means the optimal weight lies close to a quantization level.
3 Oscillations in QAT
@@
Previous studies have explored linear models to analyze the behavior of QAT and the phenomenon of weight oscillations pseudoQuantNoise; nagel2022overcoming; vitoscillations; gupta2023reducing. Inspired by these works, we also analyze a linear regression model to gain insights into the optimization dynamics during QAT.
3.1 Toy Model
@@
Consider a linear model with a single weight $w$, input $x$ and target $y$ $\in\mathbb{R}$. The quantized version of this model is defined as $\hat{y}=q(w)x$, where $q(\cdot)$ is the quantizer from Eq. (1). The quadratic loss for the quantized model is given by
$\mathcal{L}(q(w))=\frac{1}{2}(\hat{y}-y)^{2}=\frac{1}{2}(q(w)x-y)^{2}.$ (6)
@@
Our goal in this section is to understand how QAT affects the full-precision optimization process. For a given loss function $\mathcal{L}(\cdot)$ with quantized weights, we have
$\mathcal{L}(q(w))$ $=\mathcal{L}(w)+\mathcal{L}(q(w))-\mathcal{L}(w).$ (7)
We can then expand the difference in loss caused by quantization as follows
$\delta_{\text{$\mathcal{L}$}}$ $=\mathcal{L}(q(w))-\mathcal{L}(w)=\frac{1}{2}\left((q(w)x-y)^{2}-(wx-y)^{2}\right)$ (8)
$=\frac{1}{2}\left(x^{2}\left(q(w)^{2}-w^{2}\right)\right)+\left(yx(w-q(w))\right).$ (9)
@@
This expression decomposes the loss difference $\delta_{\text{$\mathcal{L}$}}$ into a quadratic term $\frac{1}{2}x^{2}(q(w)^{2}-w^{2})$ and a linear term yx(wq(w))yx(w-q(w)).
Next we derive the gradient of $\delta_{\mathcal{L}}$ wrt. $w$
$\frac{\partial\delta_{\mathcal{L}}}{\partial w}$ $=\frac{\partial}{\partial w}\bigg(\mathcal{L}(q(w))-\mathcal{L}(w)\bigg)=\frac{\partial}{\partial w}\left(\frac{1}{2}x^{2}(q(w)^{2}-w^{2})+yx(w-q(w))\right)$ (10)
$=x^{2}\left(q(w)\frac{\partial q(w)}{\partial w}-w\right)+yx\left(1-\frac{\partial q(w)}{\partial w}\right).$ (11)
@@
Using the STE and recalling that $\frac{\hat{\partial}q}{\hat{\partial}w}=1$ the expression of the STE gradient simplifies to [^2] Note that there is no clamping in the quantizer because of the scale factor Eq. (2).
$\frac{\hat{\partial}\delta_{\mathcal{L}}}{\hat{\partial}w}=x^{2}(q(w)-w)=-x^{2}\mathbf{\Delta}(w).$ (12)
3.2 Oscillation Mechanism